# Onsager-Machlup Functional and Large Deviation Principle for Stochastic Hamiltonian Systems

### v1

@@ 2.4. Technical Lemmas
The following two theorems are fundamental parts of calculating Onsager-Machup functional. Their proofs can be found in [23].
Lemma 2.8.
Let $\mathcal{N}$ be a measurable norm on $H^{1}$. For any $f\in L^{2}([0,1])$, we have
$\lim\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{{\int_{0}^{1}{f(s)}\,{\rm d}W_{s}}\right\}\big{|}\tilde{\mathcal{N}}<\epsilon\right)=1,$
where $\tilde{\mathcal{N}}$ is defined by Definition 1.
@@
We say that an operator $S:\mathbb{H}\to\mathbb{H}$ is nuclear, if
$\sum_{n=1}^{\infty}\left|\langle Se_{n},k_{n}\rangle\right|<\infty,$
for any orthonormal sequences $B_{1}=\left\{e_{n}\right\}_{n\in\mathbb{N}}$ and $B_{2}=\left\{k_{n}\right\}_{n\in\mathbb{N}}$ in $\mathbb{H}$.
We define the trace of a nuclear operator $S$ as follows:
$Tr.S=\sum_{n=1}^{\infty}\langle Se_{n},e_{n}\rangle$
@@
for any orthonormal sequence $B=\left\{e_{n}\right\}_{n\in\mathbb{N}}$ in $\mathbb{H}$. The definition of trace is independent of the orthonormal sequences we choose. For a given symmetric function $f\in L^{2}([0,1]^{2})$, the Hilbert-Schmidt operator $S(f):\mathbb{H}\to\mathbb{H}$ defined by:
$\left(S\left(f\right)\right)(h)(t)=\int_{0}^{t}{f(t,s)h(s)}\,{\rm d}s$
@@
is nuclear if $\sum_{n=1}^{\infty}\langle Se_{n},e_{n}\rangle<\infty$ for any orthonormal sequence $B=(e_{n})_{n}$ in $\mathbb{H}$. When the function $f$ is continuous and the operator $S(f)$ is nuclear, the trace of $f$ has the following expression(see [2]):
$Tr.f:=Tr.S(f)=\int_{0}^{1}{f(t,t)}\,{\rm d}t.$
@@
Furthermore, when ${f(s,t)}$ is a continuous $n\times n$ covariance kernel in the square $0\leq s,t\leq 1$, the corresponding operator $S$ is nuclear and the expression for its trace is as follows:
$Tr.f=Tr.S(f)=\int_{0}^{1}{Tr.f(t,t)}\,{\rm d}t.$
Lemma 2.10.
Let $f$ be a symmetric function in $L^{2}([0,1]^{2})$ and let $\mathcal{N}$ be a measurable norm. If $S(f)$ is nuclear, then
@@
$\lim\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{\int_{0}^{1}{\int_{0}^{1}{f(s,t)}\,{\rm d}W_{s}}\,{\rm d}W_{t}\right\}\big{|}\tilde{\mathcal{N}}<\epsilon\right)=e^{-Tr.(f)}.$
The following theorem and lemma are about the probability estimation of Brownian motion balls, which are the basis for the theorem in this article. The proof of the theorem can be found in [18].
Lemma 2.11.
Let $\left\{W(t):t\leq 0\right\}$ be a sample continuous Brownian motion in $\mathbb{R}$ and set
$\Phi_{\alpha}(\epsilon)={\rm log}\mathbb{P}\left(\|W\|_{\alpha}\leq\epsilon\right).$
If $0<\alpha<\frac{1}{2}$, then
$\lim\limits_{\epsilon\to 0}\epsilon^{\frac{2}{1-2\alpha}}\Phi_{\alpha}(\epsilon)=-C_{\alpha}$
exists with
$2^{-\frac{2\left(1-\alpha\right)}{\left(1-2\alpha\right)}}\Lambda_{\alpha}\leq C_{\alpha}\leq\left(2^{-\frac{1}{2}}\left(2^{\alpha}-1\right)\left(2^{1-\alpha}-1\right)\right)^{-\frac{2(1-\alpha)}{(1-2\alpha)}}\Lambda_{\alpha},$
where
$\Lambda_{\alpha}=\left(\frac{2}{\pi}\right)^{\frac{1}{2}}\int_{0}^{\infty}{\frac{x^{\frac{2}{1-2\alpha}}e^{-\frac{x^{2}}{2}}}{1-G(x)}}\,{\rm d}x\qquad\text{and}\qquad G(x)=\left(\frac{2}{\pi}\right)^{\frac{1}{2}}\int_{x}^{\infty}{e^{-\frac{y^{2}}{2}}}\,{\rm d}y.$
Lemma 2.12.
Let $\sigma(t)\in C([0,T],\mathbb{R}^{n\times n})$ be a diagonal matrix, and assume that there exist positive constants $M$ and $m$ such that its diagonal elements satisfy $m\leq\sigma_{i}(t)\leq M$, for $1\leq i\leq n$. We have
$\mathbb{P}\left(\|W\|_{\alpha}\leq\frac{\epsilon}{M}\right)\leq\mathbb{P}\left(\left\|\int_{0}^{t}\left\langle\sigma(t),\,{\rm d}W_{s}\right\rangle\right\|_{\alpha}\leq\epsilon\right)\leq\mathbb{P}\left(\|W\|_{\alpha}\leq\frac{\epsilon}{m}\right)$
for any $0\leq t\leq 1$. According to Lemma 2.11, we have
$\lim\limits_{\epsilon\to 0}\mathbb{P}\left(\left\|\int_{0}^{t}\left\langle\sigma(t),\,{\rm d}W_{s}\right\rangle\right\|_{\alpha}\leq\epsilon\right)\geq\lim\limits_{\epsilon\to 0}\mathbb{P}\left(\|W\|_{\alpha}\leq\frac{\epsilon}{M}\right)\geq e^{-c\left(\frac{\epsilon}{M}\right)^{-\frac{2}{1-2\alpha}}},$
where $c=\left(2^{-\frac{1}{2}}\left(2^{\alpha}-1\right)\left(2^{1-\alpha}-1\right)\right)^{-\frac{2(1-\alpha)}{(1-2\alpha)}}\Lambda_{\alpha}$.
We define the following norms on $\mathbb{H}^{1}$, respectively:
$\mathcal{N}_{g,0}(h):=\sup\limits_{t\in[0,1]}\left|\int_{0}^{t}{g(s)h^{\prime}(s)}\,{\rm d}s\right|,$
$\mathcal{N}_{g,\alpha}(h):=\sup\limits_{t\in[0,1]}\frac{|\int_{0}^{t}{g(s)h^{\prime}(s)}\,{\rm d}s-\int_{0}^{r}{g(s)h^{\prime}(s)}\,{\rm d}s|}{|t-r|^{\alpha}},\quad 0<\alpha<\frac{1}{4},$
$\mathcal{N}_{g}(h):=\mathcal{N}_{g,0}(h)+\mathcal{N}_{g,\alpha}(h),\quad 0<\alpha<\frac{1}{4}.$
In order to apply the above theorems in this paper, we need the following lemma.
Lemma 2.13.
$\mathcal{N}_{g}$ with $0<\alpha<\frac{1}{4}$ are measurable norms and we have $\tilde{\mathcal{N}}_{g}=\|\int_{0}^{1}{g(t)}\,{\rm d}W_{t}\|_{\alpha}$.
Proof.
According to the properties of norm and semi-norm, it suffices to show that $\mathcal{N}_{g,0}$ is a measurable norm and $\mathcal{N}_{g,\alpha}(h)$ is a measurable semi-norm. Below, we will only prove that $\mathcal{N}_{g,0}$ is a measurable norm, because the proof of $\mathcal{N}_{g,\alpha}$ is similar to $\mathcal{N}_{g,0}$. Fix $t\in[0,1]$ and define the continuous linear functional $\varphi_{t}:\mathbb{H}^{1}\to\mathbb{R}$ as follows:
$\varphi_{t}(h)=\int_{0}^{1}{g(s)h^{\prime}(s)}\,{\rm d}s.$
Then we can show that $|\varphi_{t}(\cdot)|$ represents a measurable norm. Define the measurable norms $\mathcal{N}_{n}(h)=\sup\limits_{0\leq j\leq 2^{n}}|\varphi_{j2^{-n}}(h)|$. In addition, we have the following convergence regarding limits $n\to\infty$:
$\tilde{\mathcal{N}}_{n}=\sup\limits_{0\leq j\leq 2^{n}}\bigg{|}\int_{0}^{j2^{-n}}{g(t)}\,{\rm d}W_{t}\bigg{|}\stackrel{{\scriptstyle\mathbb{P}}}{{\longrightarrow}}\sup\limits_{0\leq j\leq 1}\bigg{|}\int_{0}^{1}{g(t)}\,{\rm d}W_{t}\bigg{|},$
and by lemma 2.11, we have
$\mathbb{P}\bigg{(}\sup\limits_{0\leq j\leq 1}\bigg{|}\int_{0}^{1}{g(t)}\,{\rm d}W_{t}\bigg{|}\leq\epsilon\bigg{)}>0.$
According to Lemma 2.3, $\mathcal{N}_{g,0}=\lim\limits_{n\to\infty}\mathcal{N}_{n}(\cdot)$ is a measurable norm. Similarly, it is straightforward to obtain that $\mathcal{N}_{g,\alpha}(h)$ is a measurable semi-norm. Therefore, $\mathcal{N}_{g}=\lim\limits_{n\rightarrow\infty}\mathcal{N}_{n}(\cdot)$ is a measurable norm and we have $\tilde{\mathcal{N}}_{g}=\|\int_{0}^{1}{g(t)}\,{\rm d}W_{t}\|_{\alpha}$. ∎
@@
Define the parameters:
(2.2) $\theta$ $:=\mathrm{T}\mathrm{K},$ $\lambda$ $:=\log\rho^{-1},$ $\kappa$ $:=6\sigma^{-1}\lambda,$
$\tilde{r}$ $\leq\frac{r}{32d\theta},$ $\overline{r}$ $\leq\min\left\{\frac{\alpha}{2d\mathrm{K}\kappa^{\nu}},\tilde{r}\right\},$ $\breve{r}$ $:=\frac{\tilde{r}\bar{\sigma}}{16d\theta},$
$\overline{s}$ $:=s-\frac{2}{3}\sigma,$ $s^{\prime}$ $:=s-\sigma,$ $\mathrm{L}$ $:=C_{0}\frac{\theta^{2}\varepsilon}{r\breve{r}}.$
@@
(2.3) $\sigma^{-\nu}\frac{\varepsilon}{\alpha r}\leq\rho\leq\frac{1}{4},\qquad r\leq\frac{\alpha}{\mathrm{K}}\sigma^{\nu},\qquad\mathrm{L}\leq\frac{\bar{\sigma}}{3}.$
@@
(2.4) $\begin{cases}H\circ\phi^{\prime}=H^{\prime}:=K^{\prime}+P^{\prime},\\ \partial_{I^{\prime}}K^{\prime}\circ G=\partial_{I}K,\quad\det\partial_{I^{\prime}}^{2}K^{\prime}\circ G\neq 0\quad\text{on }\mathscr{D},\end{cases}$
@@
(2.5) $\|\partial_{I^{\prime}}^{2}\widetilde{K}\|_{\overline{r}/2,\mathscr{D}^{\prime}}\leq\mathrm{K}\mathrm{L},\qquad\|G-\text{id}\|_{\overline{r},\mathscr{D}}\leq\overline{r}\mathrm{L},\qquad\|\widetilde{T}\|_{\mathscr{D}^{\prime}}\leq\mathrm{T}\mathrm{L},$
@@
For the approximation of smooth functions using real-analytic functions and the uniform convergence of sequences of real-analytic functions, we refer to the relevant results in [38].
Lemma 2.15 (Jackson, Moser, Zehnder).
@@
Lemma 2.16 (Bernstein, Moser).
@@
where $s_{j}=s_{0}\kappa^{j}$ for $s_{0}>0$, $10<\kappa<$ 1, and $l\in\mathbb{R}^{+}\setminus\mathbb{Z}$. Suppose the sequence satisfies
@@
3. Onsager-Machlup functional for stochastic Hamiltonian systems

### v2

@@ 2.4 Technical theorems
The following theorem originates from the seminal work of Shepp and Zeitouni [41], where they investigated the conditional behavior of exponential moments of Itô integrals under small-ball constraints (see page 2 of their paper). In particular, they established that for any deterministic integrand in $L^{2}[0,1]$, the conditional exponential moment approaches unity as the Brownian path is conditioned to remain within an increasingly small neighborhood. We formally restate their result below:
Theorem 2.11.
Let $f\in L^{2}[0,1]$ be a deterministic function. Then, for any “reasonable” norm $\|\cdot\|$ on $C_{0}[0,1]$ (for example, the maximum norm or the Hölder norm), we have
$\lim\limits_{\epsilon\to 0}\mathbb{E}\left(\exp\left\{\int_{0}^{1}f(s)\,\mathrm{d}W_{s}\right\}\,\middle|\,\|W\|<\varepsilon\right)=1,$
where $W_{t}$ denotes a standard Brownian motion.
To extend this framework to a more general class of norms and vector-valued integrands, we adopt the concept of measurable norms on the Cameron–Martin space $\mathbb{H}^{1}$. The following two theorems, stated as Theorems 8 and 10 in [21] and due to Gilles, provide such a generalization.
Theorem 2.12.
Let $\mathcal{N}$ be a measurable norm on $\mathbb{H}^{1}$. For any deterministic function $f\in L^{2}([0,1],\mathbb{R}^{n})$, we have
$\lim\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{{\int_{0}^{1}\left\langle{f(s)},\,{\rm d}W_{s}\right\rangle}\right\}\big{|}\tilde{\mathcal{N}}\leq\epsilon\right)=1,$
where $W_{t}$ denotes an $n$-dimensional standard Brownian motion with independent components and $\tilde{\mathcal{N}}$ is defined by Definition 2.2.
@@
We say that an operator $\mathcal{S}:\mathbb{H}\to\mathbb{H}$ is nuclear, if
$\sum_{n=1}^{\infty}\left|\langle\mathcal{S}e_{n},k_{n}\rangle\right|<\infty$
for any orthonormal sequences $B_{1}=(e_{n})_{n}$ and $B_{2}=(k_{n})_{n}$ in $\mathbb{H}$.
We define the trace of a nuclear operator $\mathcal{S}$ as follows:
$\mathrm{Tr}(\mathcal{S})=\sum_{n=1}^{\infty}\langle\mathcal{S}e_{n},e_{n}\rangle$
@@
for any orthonormal sequence $B=(e_{n})_{n}$ in $\mathbb{H}$. The definition of trace is independent of the orthonormal sequences we choose. For a given symmetric function $f\in L^{2}([0,1]^{2})$, the Hilbert-Schmidt operator $\mathcal{S}(f):\mathbb{H}\to\mathbb{H}$ is defined as
$\left(\mathcal{S}\left(f\right)\right)(h)(t)=\int_{0}^{t}{f(t,s)h(s)}\,{\rm d}s$
@@
is nuclear if $\sum_{n=1}^{\infty}\langle\mathcal{S}e_{n},e_{n}\rangle<\infty$ for any orthonormal sequence $B=(e_{n})_{n}$ in $\mathbb{H}$. When the function $f$ is continuous and the operator $\mathcal{S}(f)$ is nuclear, the trace of $f$ has the following expression(see [2]):
$\mathrm{Tr}\left(f\right):=\mathrm{Tr}\left(\mathcal{S}(f)\right)=\int_{0}^{1}{f(s,s)}\,{\rm d}s.$
@@
Furthermore, when f(s,t){f(s,t)} is a continuous $n\times n$ covariance kernel in the square $0\leq s,t\leq 1$, the corresponding operator $\mathcal{S}$ is nuclear and the expression for its trace is as follows:
$\mathrm{Tr}\left(f\right)=\mathrm{Tr}\left(\mathcal{S}(f)\right)=\int_{0}^{1}{\mathrm{Tr}\left(f(s,s)\right)}\,{\rm d}s.$
Theorem 2.14.
Let $f$ be a symmetric function in $L^{2}([0,1]^{2})$ and let $\mathcal{N}$ be a measurable norm. If $\mathcal{S}(f)$ is nuclear, then
@@
$\lim\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{\int_{0}^{1}{\int_{0}^{1}{f(s,t)}\,{\rm d}W_{s}}\,{\rm d}W_{t}\right\}\big{|}\tilde{\mathcal{N}}\leq\epsilon\right)=e^{-\mathrm{Tr}(f)}.$
The following theorem provides a classical estimate for the small ball probability of Brownian motion, which serves as a foundational result for the developments in this paper. A detailed explanation can be found on page 2 of [29].
Theorem 2.15.
Let $W(t)$ be a standard Brownian motion on $\mathbb{R}$, and define the first exit time
$\tau:=\inf\{s>0:|W(s)|\geq 1\}.$
Then for any $\varepsilon>0$, the small ball probability under the supremum norm satisfies the following scaling identity:
$\mathbb{P}\left(\sup_{0\leq t\leq 1}|W(t)|\leq\varepsilon\right)=\mathbb{P}\left(\sup_{0\leq t\leq\varepsilon^{-2}}|W(t)|\leq 1\right)=\mathbb{P}\left(\tau\geq\varepsilon^{-2}\right).$
Moreover, as $\varepsilon\to 0$, the probability exhibits the following logarithmic asymptotic behavior:
$\log\mathbb{P}\left(\sup_{0\leq t\leq 1}|W(t)|\leq\varepsilon\right)\sim-\frac{\pi^{2}}{8\varepsilon^{2}}.$
Remark 1.
heorem 2.15 provides a classical small ball estimate for one-dimensional Brownian motion. Similarly, by applying the multiplication principle for independent events, an analogous small ball estimate can be derived for $n$-dimensional Brownian motion. Although the constant may depend on the dimension $n$, since $n$ is finite, this dependence can be absorbed into the constant $C$, yielding the same asymptotic rate.
@@
Define the parameters
$\theta$ $=\mathrm{T}\mathrm{K},$ $\lambda$ $=\log\rho^{-1},$ $\kappa$ $=6\sigma^{-1}\lambda,$ (4)
$\tilde{r}$ $\leq\frac{r}{32d\theta},$ $\overline{r}$ $\leq\min\left\{\frac{\alpha}{2d\mathrm{K}\kappa^{\nu}},\tilde{r}\right\}$, $\breve{r}$ $=\frac{\tilde{r}\bar{\sigma}}{16d\theta},$
$\overline{s}$ $=s-\frac{2}{3}\sigma,$ $s^{\prime}$ $=s-\sigma,$ $\mathrm{L}$ $=C_{0}\frac{\theta^{2}\varepsilon}{r\breve{r}}.$
@@
$\sigma^{-\nu}\frac{\varepsilon}{\alpha r}\leq\rho\leq\frac{1}{4},\qquad r\leq\frac{\alpha}{\mathrm{K}}\sigma^{\nu},\qquad\mathrm{L}\leq\frac{\bar{\sigma}}{3}.$ (5)
@@
$\begin{cases}H\circ\phi^{\prime}=H^{\prime}:=K^{\prime}+P^{\prime},\\ \partial_{I^{\prime}}K^{\prime}\circ G=\partial_{I}K,\quad\det\partial_{I^{\prime}}^{2}K^{\prime}\circ G\neq 0\quad\text{on }\mathscr{D},\end{cases}$ (6)
@@
$\|\partial_{I^{\prime}}^{2}\widetilde{K}\|_{\overline{r}/2,\mathscr{D}^{\prime}}\leq\mathrm{K}\mathrm{L},\qquad\|G-\text{id}\|_{\overline{r},\mathscr{D}}\leq\overline{r}\mathrm{L},\qquad\|\widetilde{T}\|_{\mathscr{D}^{\prime}}\leq\mathrm{T}\mathrm{L},$ (7)
@@
For the approximation of smooth functions by real-analytic functions and the uniform convergence of sequences of real-analytic functions, we refer to Lemma 3 and Lemma 4 in [39], respectively.
Theorem 2.17 (Jackson, Moser, Zehnder).
@@
Theorem 2.18 (Bernstein, Moser).
@@
where $s_{j}=s_{0}\kappa^{j}$ for $s_{0}>0$, $0<\kappa<1$, and $l\in\mathbb{R}^{+}\setminus\mathbb{Z}$. Suppose the sequence satisfies
@@
2.5 Extension of Theorems
In this section, we provide three important lemmas along with their proofs, which serve as the key technical tools for the subsequent analysis. It is important to emphasize that the Hamiltonian systems considered in this paper involve multiplicative noise. This represents a significant departure from the classical setup, where the diffusion term is typically assumed to be constant. Although we extended this to time-varying coefficient noise in [48], where the stochastic term takes the form $\int_{0}^{t}g(s)\,dW_{s}$, representing a linear transformation of Brownian motion, the corresponding solution remains a Gaussian process. In contrast, the setup considered here involves a multiplicative diffusion term, whose diffusion coefficient depends nonlinearly on the current state. This results in the non-Gaussian nature of the solution process.
This non-Gaussian property introduces significant analytical challenges and renders many classical tools from Onsager-Machlup functional theory inapplicable. Specifically, foundational results such as Theorems 2.12 and 2.14 no longer hold in this more general setting.
To overcome these difficulties, we first establish a small ball probability estimate Lemma 2.19, which is a crucial component of our analysis. Using this result, we extend Theorem 2.12 via the conditional expectation comparison principle, leading to Lemma 2.20, tailored to our current setup. Building on this and referencing the results of Gilles [21], we further derive an extended version of Theorem 2.14, formulated as Lemma 2.21. These generalizations form the core technical tools for the subsequent analysis.
Lemma 2.19.
[…]
$\lim\limits_{\epsilon\to 0}\mathbb{P}\left(\|\int_{0}^{1}{g(s,X_{s})}\,{\rm d}W_{s}\|\leq\epsilon\right)\geq e^{-C\epsilon^{-2}},$
where $C>0$ is a constant depending on the dimension.
Proof.
Although the integrand $g(t,X_{t})$ is stochastic and the stochastic integral $\int_{0}^{1}g(s,X_{s})\,\mathrm{d}W_{s}$ is, in general, a nonlinear functional of the Brownian motion and hence not Gaussian distributed, a lower bound for its small-ball probability can still be established under appropriate conditions. Specifically, if $g(t,X_{t})$ satisfies the uniform ellipticity condition and is twice continuously differentiable with bounded derivatives, then the desired estimate follows.
We consider the stochastic process defined by
$A(t):=\int_{0}^{t}g(s,X_{s})\,\mathrm{d}W_{s},$
where $g(s,X_{s})$ is a predictable process with respect to the natural filtration of the Brownian motion $W_{s}$, and satisfies a uniform ellipticity condition. By Itô’s isometry, we have
$\mathbb{E}\left[|A(t)-A(s)|^{2}\right]=\mathbb{E}\left[\int_{s}^{t}\mathrm{Tr}(gg^{\top})(r,X_{r})\,\mathrm{d}r\right].$
Under the uniform ellipticity assumption, there exist constants $0<\lambda\leq\Lambda$ such that
$\lambda(t-s)\leq\mathbb{E}\left[|A(t)-A(s)|^{2}\right]\leq\Lambda(t-s),$
which implies that the trace of the quadratic variation process
$\langle A\rangle_{t}=\int_{0}^{t}gg^{\top}\,\mathrm{d}r$
has linear bounds in time.
Hence, by the Dubins–Schwarz theorem, there exists an $n$-dimensional standard Brownian motion $B_{t}$ and a predictable orthogonal process $U_{t}\in O(n)$ such that
$A(t)=U_{t}\,B_{\langle A\rangle_{t}}.$
Since $U_{t}$ is orthogonal, the Euclidean norm is preserved, i.e., $|U_{t}v|=|v|$, and therefore the supremum norm satisfies
$\|A\|:=\sup_{t\in[0,1]}|A(t)|=\sup_{t\in[0,1]}|B_{\langle A\rangle_{t}}|.$
Introducing the time change $\tau=\langle A\rangle_{t}$, and noting that $\mathrm{d}\tau/\mathrm{d}t\in[\lambda,\Lambda]$, we obtain the bounds
$\sup_{t\in[0,1]}|B_{\langle A\rangle_{t}}|\leq\sup_{s\in[0,\Lambda]}|B_{s}|.$
This yields the inclusion of events:
$\left\{\sup_{s\in[0,\Lambda]}|B_{s}|\leq\epsilon\right\}\subseteq\left\{\sup_{t\in[0,1]}\left|\int_{0}^{t}g(s,X_{s})\,\mathrm{d}W_{s}\right|\leq\epsilon\right\}\quad\text{a.s.}.$
Applying the classical small ball estimate for Brownian motion (Theorem 2.15), there exists a constant $C>0$, depending on the dimension and the upper ellipticity bound $\Lambda$, such that
$\mathbb{P}\left(\|A\|\leq\epsilon\right)\geq\mathbb{P}\left(\sup_{s\in[0,\Lambda]}|B_{s}|\leq\epsilon\right)\sim\exp\left(-\frac{C}{\epsilon^{2}}\right),\quad\text{as }\epsilon\to 0.$
Lemma 2.20.
Suppose that $X_{t}$ is the solution of the Itô stochastic differential equation (2), and that assumptions (C1) and (C2) hold. Define
$A(t):=\int_{0}^{t}g(s,X_{s})\,\mathrm{d}W_{s}.$
Let $f(t)\in L^{2}([0,1],\mathbb{R}^{n})$ be a deterministic function, and define the Itô integral
$I(f):=\int_{0}^{1}\left\langle f(s),\,\mathrm{d}W_{s}\right\rangle.$
Then for supremum norm, we have
$\lim_{\epsilon\to 0}\mathbb{E}\left[\exp\left(I(f)\right)\,\bigg{|}\,\|A\|\leq\epsilon\right]=1.$
Proof.
Since $f(t)$ is deterministic and belongs to $L^{2}([0,1])$, and $W_{t}$ is a standard Brownian motion, it follows from properties of the Itô integral that
$\mathbb{E}\left[I(f)\right]=\mathbb{E}\left[\int_{0}^{1}f(s)\,\mathrm{d}W_{s}\right]=0.$
Since the exponential function is convex, Jensen’s inequality yields
$\mathbb{E}\left(\exp\left(I(f)\right)\,\middle|\,\|A\|\leq\epsilon\right)\geq\exp\left(\mathbb{E}\left[I(f)\,\middle|\,\|A\|\leq\epsilon\right]\right)=\exp(0)=1.$
Hence we have the lower bound
$\mathbb{E}\left(\exp\left(I(f)\right)\,\middle|\,\|A\|\leq\epsilon\right)\geq 1.$
On the other hand, we employ the method of conditional expectation bounding (cf. the proof of Lemma 2.19) to relate the event $\|A\|\leq\epsilon$ to the corresponding small ball event $\|W\|\leq\epsilon^{\prime}$. Under assumption (C2), for any $t\in[0,1]$, we have
$\lambda\leq\|g(t,X_{t})\|_{HS}\leq\Lambda,$
where $\|\cdot\|_{HS}$ denotes the Hilbert-Schmidt norm of matrices. Thus, we obtain the pathwise estimate
$c_{1}\lambda\|W\|\leq\|A\|\leq c_{2}\Lambda\|W\|\quad\text{a.s.}.$
The expression $\|\cdot\|$ denotes the supremum norm, and $c_{1}$ and $c_{2}$ are two constants. Therefore, the event
$\left\{\|W\|\leq\frac{\epsilon}{c_{2}\Lambda}\right\}\subset\left\{\|A\|\leq\epsilon\right\}\subset\left\{\|W\|\leq\frac{\epsilon}{c_{1}\lambda}\right\}\quad\text{a.s.}.$
In particular, we have the inequality
$\mathbb{P}\left(\|W\|\leq\frac{\epsilon}{c_{2}\Lambda}\right)\leq\mathbb{P}\left(\|A\|\leq\epsilon\right)\leq\mathbb{P}\left(\|W\|\leq\frac{\epsilon}{c_{1}\lambda}\right).$
From the small ball asymptotics in Theorem 2.11, it is known that for any deterministic function $f\in L^{2}([0,1],\mathbb{R}^{n})$, we have
$\lim_{\epsilon\to 0}\mathbb{E}\left[\exp\left(\int_{0}^{1}f(s)\,\mathrm{d}W_{s}\right)\,\middle|\,\|W\|\leq\frac{\epsilon}{c_{2}\Lambda}\right]=1,$
and
$\lim_{\epsilon\to 0}\mathbb{E}\left[\exp\left(\int_{0}^{1}f(s)\,\mathrm{d}W_{s}\right)\,\middle|\,\|W\|\leq\frac{\epsilon}{c_{1}\lambda}\right]=1.$
Hence, by the squeeze theorem, we immediately obtain
$\lim_{\epsilon\to 0}\mathbb{E}\left[\exp\left(I(f)\right)\,\middle|\,\|A\|\leq\epsilon\right]=1.$
In fact, since $g$ satisfies the uniform non-degeneracy condition, the event $\{\|A\|\leq\epsilon\}$ is asymptotically equivalent to the small ball event $\{\|W\|\leq\epsilon^{\prime}\}$ as $\epsilon\to 0$. Moreover, since the functional $\exp(I(f))$ is continuous with respect to the underlying Brownian path and equals 1 at the zero path, its conditional expectation under small ball constraints converges to 1. Therefore, by bounding the target expectation between two quantities that both converge to 1, the desired limit follows. ∎
Remark 2.
Under the same assumptions as in the previous Lemma 2.20, for any deterministic function $f(t,x)\in L^{2}([0,1],\mathbb{R}^{n})$ and any $p>1$, we have
$\lim_{\epsilon\to 0}\mathbb{E}\left[\exp\left(I(f)^{p}\right)\,\bigg{|}\,\|A\|\leq\epsilon\right]=1.$
Lemma 2.21.
Let $A(t):=\int_{0}^{t}g(s,X_{s})\,\mathrm{d}W_{s}$, where the matrix-valued function $g(s,X_{s})$ satisfies assumption (C2). Let $f\in L^{2}([0,1]^{2})$ be a symmetric function such that the associated Hilbert–Schmidt operator $K(f)$ is nuclear. Then for supremum norm, we have
$\lim_{\epsilon\to 0}\mathbb{E}\left[\exp\left\{\mathcal{S}(f)\right\}\,\middle|\,\|A\|\leq\epsilon\right]=e^{-\mathrm{Tr}(f)},$
where $\mathcal{S}(f):=2\int_{0<s<t<1}f(s,t)\,\mathrm{d}W_{s}\,\mathrm{d}W_{t}$ denotes the second-order Skorokhod integral associated with the kernel $f$.
Proof.
We begin by applying the method of conditional expectation bounding (cf. the proof of Lemma 2.19) to relate the conditioning event $\|A\|<\epsilon$ to a small-ball event of the form $\|W\|<\epsilon^{\prime}$. Under assumption (C2), we have for all $t\in[0,1]$,
$\lambda\leq\|g(t,X_{t})\|_{HS}\leq\Lambda,$
where $\|\cdot\|_{HS}$ denotes the Hilbert-Schmidt norm of a matrix. This yields the pathwise estimate
$c_{1}\lambda\|W\|\leq\|A\|\leq c_{2}\Lambda\|W\|\quad\text{a.s.},$
and the following inclusion of events:
$\left\{\|W\|\leq\frac{\epsilon}{c_{2}\Lambda}\right\}\subset\left\{\|A\|\leq\epsilon\right\}\subset\left\{\|W\|\leq\frac{\epsilon}{c_{1}\lambda}\right\}\quad\text{a.s.}.$
This leads to bounds on small-ball probabilities
$\mathbb{P}\left(\|W\|\leq\frac{\epsilon}{c_{2}\Lambda}\right)\leq\mathbb{P}\left(\|A\|\leq\epsilon\right)\leq\mathbb{P}\left(\|W\|\leq\frac{\epsilon}{c_{1}\lambda}\right).$
Recall that the second-order Skorokhod integral satisfies the identity
$\mathcal{S}(f)=I^{2}(f)-\mathrm{Tr}(f),$
where $I^{2}(f):=\int_{0}^{1}\int_{0}^{1}f(s,t)\,\mathrm{d}W_{s}\,\mathrm{d}W_{t}$ and $\mathrm{Tr}(f):=\sum_{i,j}\lambda_{i,j}\langle h_{i},h_{j}\rangle$ is the trace of the nuclear operator $K(f)$.
According to Theorem 8 in Gilles [21], we have the exponential identity
$\lim_{\epsilon\to 0}\mathbb{E}\left[\exp\left\{I^{2}(f)\right\}\,\middle|\,\|W\|\leq\frac{\epsilon}{c_{2}\Lambda}\right]=1,$
and
$\lim_{\epsilon\to 0}\mathbb{E}\left[\exp\left\{I^{2}(f)\right\}\,\middle|\,\|W\|\leq\frac{\epsilon}{c_{1}\lambda}\right]=1.$
Then, by the squeeze theorem, we deduce that
$\lim_{\epsilon\to 0}\mathbb{E}\left[\exp\left\{I^{2}(f)\right\}\,\middle|\,\|A\|\leq\epsilon\right]=1,$
which yields
$\lim_{\epsilon\to 0}\mathbb{E}\left[\exp\left\{\mathcal{S}(f)\right\}\,\middle|\,\|A\|\leq\epsilon\right]=e^{-\mathrm{Tr}(f)}.$
3 Onsager-Machlup functional for stochastic Hamiltonian systems